The pH of a 0.00156 M solution of a certain weak acid, HA is found to be 3.64. What is the Ka of acid?

2.37 x 10-4

3.94 x 10-5

2.81 x 102

4.29 x 10-4
Answer #1

HA <-------------> H+ + A-

pH = 3.64

[H+] = 10-pH = 10-3.64 = 2.30 x 10-4 M

at equilibrium [H+] = [A-] = 2.30 x 10-4 M

[HA] = 0.00156 - 0.000230 = 0.00133 M

Ka = [H+] [A-] / [HA]

Ka = [2.30 x 10-4] [2.30 x 10-4] / [0.00133]

Ka = 3.9 x 10-5

answer = option 2 = 3.94 x 10-5